# A Heterogeneous Dynamic Convolutional Neural Network for Image Super-resolution

### v2

@@ V Conclusion
In this paper, we propose a heterogeneous dynamic convolutional network in SISR. This paper designs a heterogeneous dynamic convolutional network to capture more structural information. The upper network depends on stacked heterogeneous blocks to facilitate more contexture information for improving effects of image super-resolution. Also, each heterogeneous block is composed of a dilated, dynamic, common convolutional layers, ReLU and residual learning operation is used to adjust parameters for different inputs and prevent long-term dependency problem. The lower network uses symmetric architecture to enhance relations of different layers to extract more structural information for SISR. We will deal with SISR with non-reference images in the future.

### v3

@@ V Conclusion
In this paper, we propose a adaptive convolutional network via a parallel network in SISR. The upper network depends on enhancing relation of context information, salient information relation of a kernel mapping and relations of shallow and deep layers to implement efficient SR performance. The lower network relies on a symmetric architecture to strengthen relation of hierarchical information to improve effect of SISR. Two networks can extract complementary information to facilitate more robust information to protect performance of SISR for different scenes. We will deal with SISR with non-reference images in the future.